OpenStudy (anonymous):

6-w/w-3 - 8w-7/3-w subtract 6-w/w-3 - 8w-7/3-w subtract @Mathematics

OpenStudy (anonymous):

\[-9x-(1/3)=0 , w=-1/27\]

OpenStudy (anonymous):

its w

OpenStudy (anonymous):

i get lost with the lcm

OpenStudy (anonymous):

is your question\[6-(w/w)-3-(8w)-(7/3)-w\]

OpenStudy (anonymous):

(1-7w)/(w-3)

OpenStudy (anonymous):

(6-w)/(w-3)-(8*w-7)/(3-w) = (6-w)/(w-3)-(-8*w+7)/(w-3)

OpenStudy (anonymous):

ugh ok im coming up with different answers what did yall get g=for the lcm

OpenStudy (anonymous):

g?

OpenStudy (anonymous):

i mean get,, sorry

OpenStudy (anonymous):

u write the question in the proper manner with brackets

OpenStudy (anonymous):

just multiply the numerator and denominator of the second fraction with (-1)

OpenStudy (anonymous):

then you'll get what i posted

OpenStudy (anonymous):

u will only get (1-7w)/(w-3)

OpenStudy (anonymous):

ok i see what your saying thanks.

OpenStudy (anonymous):

pleasure